Presupposition

Hans Kamp

(editor)

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852

Deliverable R2.2.A, Part II August 1993

Presupposition

Hans Kamp Universität Stuttgart

(editor)

authors:

David Beaver Kjell Johan Sæbø

Henk Zeevat Rob van der Sandt

DYANA-2

Dynamic Interpretation of Natural Language

esprit Basic Research Project 6852

Deliverable R2.2.A, Part II August 1993 esprit Basic Research Action 6852 dyana-2

Dynamic Interpretation of Natural Language

Universiteit van Amsterdam Institute for Logic, Language and Computation (ILLC)

University of Edinburgh Centre for Cognitive Science (ECCS)

Universität München Centrum für Informations- und Sprachforschung (CIS)

Universitetet i Oslo Department of Linguistics and Philosophy (ILF)

Universität Stuttgart Institut für Maschinelle Sprachverarbeitung (IMS)

Universität Tübingen Seminar für Sprachwissenschaft (SfS)

Rijksuniversiteit Utrecht Research Institute for Language and Speech (OTS)

For copies of reports, updates on project activities and other dyana-related information, contact:

The dyana-2 Project Administrator ILLC/Department of Philosophy University of Amsterdam Nieuwe Doelenstraat 15 NL-1012 CP Amsterdam

\begin{matrix}c \\ ◯ \end{matrix} 1993, The Individual Authors, except:

Presupposition and Accomodation in Update Semantics, \begin{matrix}c \\ ◯ \end{matrix} Oxford University Press

No part of this document may be reproduced or transmitted in any form, or by any means, electronic or mechanical, including photocopy, recording, or any information storage and retrieval system, without permission from the copyright owner.

Contents

Introduction by the editor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . v

Task 2.2, subtask 8: Presupposition Two Birds and One Stone David Beaver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

The Kinematics of Presupposition David Beaver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

Kinematic Montague Grammar David Beaver . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

Additional Contributions Anaphoric Presuppositions, Accomodation, and Zero Anaphora Kjell Johan Sæbø . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

Presupposition and Accomodation in Update Semantics Henk Zeevat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 193

Comments on Beaver, Sæbø and Zeevat Resolution and Accomodation Rob van der Sandt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 231

Introduction

The present deliverable contains a number of papers on the subject of presupposition, three by David Beaver, one each by Henk Zeevat and Kjell Johan Sæbø and rounded off by a commentary by Rob van der Sandt. These papers are all about presupposition. But they can fairly be said to be also about something more general – about the dynamic nature of natural language semantics and – even beyond that – about the dynamics of information. It should not surprise that presuppositional phenomena should be used as a focus for investigations with these wider aims. For as much as any other area in the study of language presupposition has, virtually from the moment it came to be perceived as an important topic for linguistics, pointed unequivocally towards the dynamic aspects of semantics (or, as one would have been more likely to put it at the time, towards the intimate connections between semantics and pragmatics.) At the very latest this perspective had become inescapable once the Projection Problem was recognized as the central challenge that any theory of presupposition had to meet and as a central challenge for any semantics and pragmatics of natural language generally. In the seventies thinking about presupposition crucially influenced our perception of the relation between semantics and pragmatics (or better: what was thought of as semantics and pragmatics at the time). In fact, some of the outstanding contributions to the study of presupppositions from that period – Karttunen, Stalnaker, Gazdar – achieved what they did achieve in large part by shedding new light on how semantics and pragmatics were connected. Since the early eighties the perspective has shifted – from what was seen as the semantics-pragmatics interface to the conception of a dynamic semantics, which incorporates not only all that had always been thought to be part of semantics, but also much that was previously considered to belong within the realm of pragmatics. The decisive paper was Heim’s terse contribution to the subject of 1983, which showed that within the dynamic semantics she was advocating the core cases of the Projection Problem simply ceased to be problems at all. Heim’s paper can now be seen as pointing clearly towards much of what the recent work on presupposition is pursuing in greater detail. Yet there has been something of a delay: the surge of interest in presuppositional phenomena is a thing of the past four years or so. Looking back one might wonder why a good part of the results that have appeared over these last four years did not happen four or five years earlier, for as Heim’s paper shows, much of the perspective and many of the tools were already in place. Be this as it may, presupposition theory has become once more one of the most active areas in the semantics of natural language. That it should be of particular importance within Dyana hardly needs comment: for a project in which the dynamic aspects of language are pivotal, presupposition is a domain of application par excellence. One of the most important recent insights about presupposition concerns its connections with anaphora. This side to the phenomenon of presupposition, which has been stressed particularly strongly in the work of van der Sandt, is particularly clearly visible in the present contribution of Sæbø, as well as in the one of Zeevat. This is the insight that there are close similarities between the contextual verification of presupposition on the one hand and the contextual resolution of an anaphoric expression (e.g. an anaphoric pronoun) on the other. As a consequence anaphora comes to be seen as often involving the “resolution” of information structures considerably more complex than the old paradigms of anaphora – pronouns, tenses, adverbs such as ‘the next day’ – and so one is led to an even more comprehensive picture, which includes besides presupposition and anaphora also ellipsis and much in the realm of focus. The complexity of these information structures and of their verification and accommodation is one aspect of recent presupposition theory which plays a major part in the contributions to this deliverable. Another aspect prominent in this deliverable – it is a central theme in the contributions by David Beaver – is one that might be said to point in the opposite direction. Verification and accommodation of presuppositions can be seen as special instances of testing and updating in dynamic theories of information such as Veltman’s Update Semantics. So it is natural to study these aspects of presupposition at a certain level of abstraction which is suited to reveal their more general information-theoretic properties. The need for studying the phenomena also at such a more abstract level has dramatically increased as the complexity and variety of presuppositional and related phenomena that are being subjected to individual scrutiny have grown – the more and the bigger the trees, and the more closely we look at their individual branching and foliage, the greater the danger of loosing sight of the wood; and the more important it is to regain a proper distance, from which the wood becomes once again surveyable. It is thus that these two aspects of the theory of presupposition, (and of dynamic semantics generally) naturally complement each other.

Hans Kamp

Task 2.2, subtask 8 Presupposition

Two Birds and One Stone The Kinematics of Presupposition Kinematic Montague Grammar

Anaphoric Presuppositions, Accomodation, and Zero Anaphora

Kjell Johan Sæbø

(University of Oslo)

Presupposition and Accomodation in Update Semantics

Henk Zeevat

(University of Amsterdam)

Comments on Beaver, Sæbø and Zeevat

Resolution and Accomodation

Rob van der Sandt

(University of Nijmegen)